OpenStudy (anonymous):

determine the derivative of the following functions x-squared over x-1

OpenStudy (anonymous):

\[\frac{d}{dx}\frac{x^2}{x-1}=\frac{x^2-2x}{(x-1)^2}\]\]

OpenStudy (anonymous):

need the steps? straight up quotient rule. use \[(\frac{f}{g})'=\frac{gf'-fg'}{g^2}\] \[f(x)=x^2,f'(x)=2x,g(x)=x-1,g'(x)=1\] and plug them in

OpenStudy (anonymous):

satellite73 derivatives have its step by step process right...

OpenStudy (anonymous):

it should be \[x^2/x-1\]

OpenStudy (anonymous):

only work is \[g'f-f'g=(x-1)2x-x^2\times 1=x^2-2x\]\]

OpenStudy (anonymous):

that is the only "step" you need really. just the algebra in the numerator

OpenStudy (anonymous):

then \[x \over (x+Deltax)-1\]

OpenStudy (anonymous):

together with \[x \over x-1\]